OpenStudy (anonymous):

Simplify each expression. Use positive exponents. M^3 n^-6 p^0 show work

zepdrix (zepdrix):

\[\large\rm m^3\color{orangered}{n^{-6}}\color{royalblue}{p^0}\]To deal with the orange part, we'll use this rule:\[\large\rm \color{orangered}{x^{-a}=\frac{1}{x^a}}\]We flip it! And the exponent becomes positive. Do deal with the blue part, Anything to the 0 power is 1. You can think of it like... p^1 is p. So p^0 is dividing p by itself.

zepdrix (zepdrix):

\[\large\rm m^3\color{orangered}{n^{-6}}\color{royalblue}{p^0}=m^3\color{orangered}{n^{-6}}\color{royalblue}{1}\]So that rule takes care of the blue part. Do you understand how the orange rule helps us? :D

OpenStudy (anonymous):

I'm lost.. I seriously suck at algebra. I've been stuck on this question for 7 months

zepdrix (zepdrix):

7 months?? 0_o oh boy

zepdrix (zepdrix):

Anyway, exponent rule tells us to flip it, and change the exponent to positive,\[\large\rm m^3\color{orangered}{n^{-6}}\color{royalblue}{p^0}=m^3\color{orangered}{\frac{1}{n^6}}\color{royalblue}{1}\]

OpenStudy (anonymous):

Hold up. M^3 N^-6 P^0 M^3 x 1/n^6 =m^3/n^6

OpenStudy (anonymous):

What now o.o

zepdrix (zepdrix):

nothing else :o

OpenStudy (anonymous):

You mean.... I'm finally done O.O

zepdrix (zepdrix):

lol ya :U